# An Actionable Framework for Assessing Bias and Fairness in Large Language Model Use Cases

### v1

@@ 3.2 Fairness Metrics for Classification Use Cases
For the remainder of this section, assume each prompt in $\mathcal{P}_{X}$ for a given classification LLM use case corresponds to a protected attribute group. Under this assumption, traditional machine learning fairness metrics (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023) may be applied (Czarnowska et al., 2021). Accordingly, these metrics are defined on binary predictions $\hat{Y}_{1},...,\hat{Y}_{N}$, generated from a sample of prompts $X_{1},...,X_{N}\in\mathcal{P}_{X}$, with some metrics also incorporating corresponding ground truth values $Y_{1},...,Y_{N}$. These metrics effectively assess group fairness (see Section 2.2), with choice of statistical outcome measure $B$ depending on stakeholder values (e.g. the relative cost of false negatives vs. false positives).
@@ 3.2.1 Representation Fairness Metrics for Binary Classification
Representation fairness metrics aim to determine whether protected attribute groups are adequately represented in the positive predictions generated by classifier. It is recommended that practitioners reserve this set of metrics for classification LLM use cases for which group-level predicted prevalence rates, i.e. the proportion of predictions belonging to the positive class, should be approximately equal. [^21] For example, when predicting which applicants are qualified for a job it may be desirable to target equal rates of positive predictions for males and females. However, when predicting certain diseases, this may not be a desirable model behavior. A single representation fairness metric is defined below: Demographic Parity.
@@
where $\hat{Y}$ denotes a model prediction and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample prompts drawn from $\mathcal{P}_{X}$.
@@ 3.2.2 Error-Based Fairness Metrics for Binary Classification
Error-based fairness metrics aim to determine whether disparities in model performance exist across protected attribute groups. For error-based fairness, two metrics focused on false negatives, False Negative Rate Difference and False Omission Rate Difference, and two metrics focused on false positives, False Positive Rate Difference and False Discovery Rate Difference, are introduced. Following Saleiro et al. (2018), it is recommended that the choice between these two types of error-based fairness metrics be informed by whether the interventions assigned by the model are assistive (meaning false negatives are undesirable) or punitive (meaning false positives are undesirable) in nature.
@@
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$. Note that false negative rate measures the proportion of actual positives $(Y=1)$ that are falsely classified as negative $(\hat{Y}=0)$. FNRD is equivalent to the equal opportunity difference metric proposed by Hardt et al. (2016).
@@
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based predictions generated from on a sample of prompts drawn from $\mathcal{P}_{X}$. Instead of concentrating on actual positives, false omission rate calculates the percentage of predicted negatives ($\hat{Y}=0$) that are misclassified. Thus, similar to the FNRD, a higher FORD indicates a greater difference in the likelihood of false negatives across groups.
@@
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$. Note that false positive rate measures the percentage of actual negatives ($Y=0$) being incorrectly predicted as positive $(\hat{Y}=1)$.
@@
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$. Rather than considering actual negatives, false discovery rate calculates the proportion of predicted positives ($\hat{Y}=1$) that are incorrectly classified. Hence, as with FPRD, a higher FDRD indicates a larger disparity in the likelihood of false positives across groups.
@@ 3.3 Fairness Metrics for Recommendation Use Cases
This section outlines a set of fairness metrics for recommendation LLM use cases, as proposed by Zhang et al. (2023). To ensure consistency with the metrics discussed in Section 3.1.3, this section presents modified versions of these metrics to be pairwise in nature, rather than attribute-wise. Given two protected attribute groups $G^{\prime},G^{\prime\prime},$ and an LLM use case $(\mathcal{M}^{(R)},\mathcal{P}_{X})$, these metrics assess similarity in counterfactually generated recommendation lists. Below, each metric is defined according to responses generated from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ which are drawn from $\mathcal{P}_{X|\mathcal{A}}$. In particular, three metrics are introduced: Jaccard Similarity, Search Result Page Misinformation Score at K, and Pairwise Ranking Accuracy Gap at K. Each of these metrics ranges in value from 0 to 1, with larger values indicating a greater degree of fairness.
@@ 4 A Unified Framework for Bias and Fairness Assessments of LLM Use Cases
In defining the proposed decision framework for selecting LLM bias and fairness evaluation metrics, the scope is restricted to use cases for which prompts can be sampled from a known population and the task is well-defined. Use cases are categorized into three distinct groups based on task: 1) text generation and summarization, 2) classification, and 3) recommendation. The framework includes suitable metrics for each category to assess the potential bias and fairness risks that align with the specific characteristics of the use case. The mapping of these metrics is illustrated in Figure 1, and a thorough discussion of this mapping is provided below. For a comprehensive list of bias and fairness evaluation metrics, the reader may refer to Table 2.
@@
Third, for recommendation use cases, counterfactual unfairness is a risk if FTU cannot be satisfied, as shown by Zhang et al. (2023). Note that counterfactual invariance may not be a desirable property for certain recommendation use cases. For instance, it may be preferred to recommend different products for male vs. female customers. Hence, if counterfactual invariance is a desired property, it is recommended that recommendation use cases not satisfying FTU be assessed for counterfactual unfairness in recommendations using the metrics outlined in 3.3. On the other hand, if a recommendation use case satisfies FTU or if counterfactual invariance is not desired, then a fairness assessment is not applicable.

### v2

@@ 3.2 Fairness Metrics for Classification Use Cases
For the remainder of this section, assume each prompt in $\mathcal{P}_{X}$ for a given classification LLM use case corresponds to a protected attribute group. Under this assumption, traditional machine learning fairness metrics (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023) may be applied (Czarnowska et al., 2021). Accordingly, these metrics are defined on binary predictions $\hat{Y}_{1},...,\hat{Y}_{N}$, generated from a sample of prompts $X_{1},...,X_{N}\in\mathcal{P}_{X}$, with some metrics also incorporating corresponding ground truth values $Y_{1},...,Y_{N}$. These metrics effectively assess group fairness (see Section 2.2) between two protected attribute groups $G^{\prime}$ and $G^{\prime\prime}$, with choice of statistical outcome measure $B$ depending on stakeholder values (e.g. the relative cost of false negatives vs. false positives).
@@ 3.2.1 Representation Fairness Metrics for Binary Classification
Representation fairness metrics aim to determine whether protected attribute groups are adequately represented in the positive predictions generated by classifier. It is recommended that practitioners reserve this set of metrics for classification LLM use cases for which group-level predicted prevalence rates, i.e. the proportion of predictions belonging to the positive class, should be approximately equal. [^21] For example, when predicting which applicants are qualified for a job it may be desirable to target equal rates of positive predictions for males and females. However, when predicting certain diseases, this may not be a desirable model behavior. A single representation fairness metric, known as Demographic Parity, is defined below.
@@
where $\hat{Y}$ denotes a model prediction and $P(\cdot)$ represents the empirical probability based on predictions generated from a sample prompts drawn from $\mathcal{P}_{X}$.
@@ 3.2.2 Error-Based Fairness Metrics for Binary Classification
Error-based fairness metrics aim to determine whether disparities in model performance exist across protected attribute groups. For error-based fairness, two metrics focused on false negatives, False Negative Rate Difference and False Omission Rate Difference, and two metrics focused on false positives, False Positive Rate Difference and False Discovery Rate Difference, are presented. Following Saleiro et al. (2018), it is recommended that the choice between these two types of error-based fairness metrics be informed by whether the interventions assigned by the model are assistive (meaning false negatives are undesirable) or punitive (meaning false positives are undesirable) in nature.
@@
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ represents the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$. Note that false negative rate measures the proportion of actual positives $(Y=1)$ that are falsely classified as negative $(\hat{Y}=0)$. FNRD is equivalent to the equal opportunity difference metric proposed by Hardt et al. (2016).
@@
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ represents the empirical probability based predictions generated from on a sample of prompts drawn from $\mathcal{P}_{X}$. Instead of concentrating on actual positives, false omission rate calculates the percentage of predicted negatives ($\hat{Y}=0$) that are misclassified. Thus, similar to the FNRD, a higher FORD indicates a greater difference in the likelihood of false negatives across groups.
@@
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ represents the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$. Note that false positive rate measures the percentage of actual negatives ($Y=0$) being incorrectly predicted as positive $(\hat{Y}=1)$.
@@
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ represents the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$. Rather than considering actual negatives, false discovery rate calculates the proportion of predicted positives ($\hat{Y}=1$) that are incorrectly classified. Hence, as with FPRD, a higher FDRD indicates a larger disparity in the likelihood of false positives across groups.
@@ 3.3 Fairness Metrics for Recommendation Use Cases
This section outlines a set of fairness metrics for recommendation LLM use cases, as proposed by Zhang et al. (2023). To ensure consistency with the metrics discussed in Section 3.1.3, this section presents modified versions of these metrics to be pairwise in nature, rather than attribute-wise. Given two protected attribute groups $G^{\prime},G^{\prime\prime},$ and an LLM use case $(\mathcal{M}^{(R)},\mathcal{P}_{X})$, these metrics assess similarity in counterfactually generated recommendation lists. Below, each metric is defined according to responses generated from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ which are drawn from $\mathcal{P}_{X|\mathcal{A}}$. In particular, three metrics are presented: Jaccard Similarity, Search Result Page Misinformation Score at K, and Pairwise Ranking Accuracy Gap at K. Each of these metrics ranges in value from 0 to 1, with larger values indicating a greater degree of fairness.
@@ 4 A Unified Framework for Bias and Fairness Assessments of LLM Use Cases
The scope of this framework is restricted to use cases for which prompts can be sampled from a known population and the task is well-defined. Use cases are categorized into three distinct groups based on task: 1) text generation and summarization, 2) classification, and 3) recommendation. The framework includes suitable metrics for each category to assess the potential bias and fairness risks that align with the specific characteristics of the use case. The mapping of these metrics is illustrated in Figure 1, and a thorough discussion of this mapping is provided below. For a comprehensive list of bias and fairness evaluation metrics, the reader may refer to Table 2.
@@
Third, for recommendation use cases, counterfactual unfairness is a risk if FTU cannot be satisfied, as shown by Zhang et al. (2023). Note that counterfactual invariance may not be a desirable property for certain recommendation use cases. For instance, it may be preferred to recommend different products for male vs. female customers. Hence, if counterfactual invariance is a desired property, it is recommended that recommendation use cases not satisfying FTU be assessed for counterfactual unfairness in recommendations using the metrics outlined in Section 3.3. On the other hand, if a recommendation use case satisfies FTU or if counterfactual invariance is not desired, then a fairness assessment is not applicable.